OpenStudy (anonymous):

Apply exponent : (-3w^7m^7)^3

OpenStudy (skullpatrol):

Use the rule of exponents for a power of a power.

OpenStudy (skullpatrol):

Or write (-3w^7m^7)^3 as (-3w^7m^7)(-3w^7m^7)(-3w^7m^7)

OpenStudy (anonymous):

(-3w^21m^21)

OpenStudy (skullpatrol):

You missed the -3?

OpenStudy (skullpatrol):

(-3)(-3)(-3)

OpenStudy (anonymous):

No, look ---> (-3w^21m^21)

OpenStudy (skullpatrol):

(-3w^7m^7)(-3w^7m^7)(-3w^7m^7) =(-3)(w^7)(m^7)(-3)(w^7)(m^7)(-3)(w^7)(m^7)

OpenStudy (skullpatrol):

See (-3)(-3)(-3)?

OpenStudy (skullpatrol):

(-3w^7m^7)^3 =(-3w^7m^7)*(-3w^7m^7)*(-3w^7m^7) =(-3)(w^7)(m^7)(-3)(w^7)(m^7)(-3)(w^7)(m^7)

OpenStudy (skullpatrol):